Find the root of the equation x^3 - x - 11 = 0 using bisection method.

up to six iterations.


1
Expert's answer
2021-04-15T06:57:00-0400

"x^3 - x - 11 = 0"

"\text{let } f(x) = x^3 - x - 11"

"x= 0 ;f(x)= 0-0-11<0"

"x=1;f(x)=1-1-11<0"

"x=2;f(x)= 8-2-11<0"

"x=3;f(x)=27-3-11>0"

"x\isin[2;3]\exists x" :f(x) =0

"\text{1 th iteration}"

"[2;3] ;c=\frac{3+2}{2}=2.5"

"f(2.5)=2.5^3-2.5-11=2.125>0"

"\text{2 th iteration}"

"[2;2.5];c=\frac{2+2.5}{2}=2.25"

"f(2.25)=2.25^3-2.25-11=-1.859375<0"

"\text{3 th iteration}"

"[2.25;2.5];c=\frac{2.25+2.5}{2}=2.375"

"f(2.375)=2.375^3-2.375-11=0.021484375>0"

"\text{4 th iteration}"

"[2.25;2.375];c=\frac{2.25+2.375}{2}=2.3125"

"f(2.3125)=2.3125^3-2.3125-11=-0.946044921875<0"

"\text{5 th iteration}"

"[2.3125;2.375];c=\frac{2.3125+2.375}{2}=2.34375"

"f(2.34375)=2.34375^3-2.34375-11=\newline
=-0.469146728515625<0"

"\text{6 th iteration}"

"[2.34375;2.375];c=\frac{2.34375+2.375}{2}=2.359375"


Answer: 2.359375 the root of the equation after 6 iterations

on the interval [2;3]